Three boys agree to meet at midday tomorrow. The probability of each boy remembering

to meet is 3/4 , 4/5 , 5/8


1
Expert's answer
2020-06-14T17:41:22-0400

Consider the following 4 events:

"A_{011}" - “only 1st boy will forget to meet”

"A_{101}" - “only 2nd boy will forget to meet”

"A_{110}" - “only 3rd boy will forget to meet”

"A_{111}" - “all three boys will remember to meet”

These events are collectively exhaustive. We should compute the probability "p" that at least two of the boys will remember to meet, i.e. the probability


"p=P(A_{011}\cup A_{101}\cup A_{110}\cup A_{111})"

Since the events are collectively exhaustive it follows that


"p=P(A_{011}\cup A_{101}\cup A_{110}\cup A_{111})="

"=P(A_{011})+P( A_{101})+P( A_{110})+P( A_{111})"

"A_{011}=A_1^C\cap A_2\cap A_3"

"A_{101}=A_1\cap A_2^C\cap A_3"

"A_{110}=A_1\cap A_2\cap A_3^C"

"A_{111}=A_1\cap A_2\cap A_3"

Given "P(A_1)=\dfrac{3}{4}, P(A_2)=\dfrac{4}{5}, P(A_3)=\dfrac{5}{8}."

Assume that events "A_1, A_2" and "A_3" are independent. This means that


"P(A_{011})=P(A_1^C)P(A_2)P(A_2)="

"=(1-\dfrac{3}{4})(\dfrac{4}{5})(\dfrac{5}{8})=\dfrac{1}{8}"

"P(A_{101})=P(A_1)P(A_2^C)P(A_2)="

"=\dfrac{3}{4}(1-\dfrac{4}{5})(\dfrac{5}{8})=\dfrac{3}{32}"

"P(A_{110})=P(A_1)P(A_2)P(A_2^C)="

"=\dfrac{3}{4}(\dfrac{4}{5})(1-\dfrac{5}{8})=\dfrac{9}{40}"

"P(A_{111})=P(A_1)P(A_2)P(A_2)="

"=\dfrac{3}{4}(\dfrac{4}{5})(\dfrac{5}{8})=\dfrac{3}{8}"

Hence the probability "p" that at least two of the boys will remember to meet is equal to 


"p=P(A_{011})+P( A_{101})+P( A_{110})+P( A_{111})="

"=\dfrac{1}{8}+\dfrac{3}{32}+\dfrac{9}{40}+\dfrac{3}{8}=\dfrac{131}{160}=0.81875"